OpenStudy (anonymous):

Let f be the function defined by: f(x)= {sin x, x<0 {x^2, 02 For what values of x is f not continuous? A. 0 only B. 1 only C. 2 only (teacher said this was it) D. 0 and 2 only E. 0, 1, and 2

OpenStudy (anonymous):

It is C. To check continuity for these piecewise functions, the upper limit for the function must equal the lower limit for where the next functions starts. In English...sin (0) must equal 0 ^2, 1^2 must equal 2 - 1, and so on. Pretty much, where each part ends, the next function must pick up at the same point. The only one that doesn't is at x = 2. Because 2 - 2 = 0, but 2 - 3 = -1. Hope you understood my messy explanation.

OpenStudy (anonymous):

How would you go about doing the problem? I get overwhelmed with all of it - especially the sin part.

OpenStudy (anonymous):

@dtcrane

OpenStudy (anonymous):

Yes, it's hard to explain over the internet. Look at it this way. F is not continuous if you have to lift your pencil off the graph. For instance, y = x. You can draw this line forever, and ever, and ever, and...and never have to lift your pencil off the page. This function, you have to at x = 2. f(x)= {sin x, x<0 {x^2, 0<x<1 {2-x, 1<x<2 {x-3, x>2 Imagine drawing sin x where x < 0. You're drawing, and drawing, and as you come closer to x = 0, sin (x) = 0. Then once you get close to 0, you have to start drawing x^2. Do you have to lift the pencil off the paper? No! Because sin(x) ended near 0, and x^2 picked up right where sin(x) left off. Again, now you have to draw x^2 from 0 to 1. So you're drawing, and drawing...as you get very very close to x = 1, x^2 = 1, right? Then, you have to start drawing 2-x. Do you have to take your pencil off the graph? No again! Because x^2 at 1 = 1, and so does 2 - x! Now try it for 2 - x up to 2, and see what happens when you have to start drawing x-3. Let me know if you understand this better please!

OpenStudy (anonymous):

I don't understand this at all :3

OpenStudy (anonymous):

I understand the part about not lifting your pencil - my teacher explained that concept. But when you said "This function, you have to at x = 2" I get confused. How did you know to do that?

OpenStudy (anonymous):

I need to be able to do this algebraically - without drawing or using a calculator

OpenStudy (anonymous):

Because 2 - x is only for 1< x < 2, right? That means when x = 2, plugging in, 2 - x ends at 0. Because 2 - x = 2 - 2 = 0 at x = 2. And x - 3 is for 2 < x < 3, right? So at 2, x - 3 = 2 - 3 = -1. So the endpoints of each tiny function have to be equal for it to be continuous. Since these two aren't equal at x = 2, it is not continuous.

OpenStudy (anonymous):

How did you plug it in? - the 2-x I mean

OpenStudy (anonymous):

We know the function is 2 - x, right? So when x = 2, you can plug it in for x. so 2 - x AT x = 2, is 2 - 2 = 0.

OpenStudy (anonymous):

why did you choose to work with 2-x? what about the other three pieces?

OpenStudy (anonymous):

You have to do all of them. You have to plug in all the endpoints to make sure they are equal. we did it for x = 0 ( sin(x) and x ^2), x = 1 (x^2 and 2 - x). Those were both equal at 0 or 1, so they were continuous. I'm having a hard time trying to explain this better, so if you don't understand after this, I am very sorry!